RC meeting and Run Plan, Wednesday November 16th
From SBS wiki
Return to the RC Meeting Calendar
RC Meeting Information
Time:at 3:30pm today
- Location: Counting house conference room
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
Accelerator and Hall Status
- Beam came back around 1:30am. We were finishing up some target testing, so couldn't take beam immediately.
- Around 3am we started beam checkout. Tuning beam to the hall took a long time, as expected! At 6:30am we started taking data on H2. Thanks to Simona for being there the whole time.
- BB GEM UV layer0 HV tripped. GEM on call was contacted and the issue was fixed quickly. Thank you Manju. 4084885
- Gordon instructed to do an NMR sweep at around 9:00 am and then start production. NMR sweep at 9:20am showed 43% polarization. We are taking production since then. 4084870
- Moller measurement is scheduled for tomorrow day shift.
- HWien angle was changed from -14.2 to 0.0 deg yesterday. The laser spot was moved to a new location with higher QE as well. 4084408
- Dump right Ion chamber unresponsive. For now director of operations agreed to allow running with just one dump IC. - Dump right IC is still down! Jiwan will follow up with Jerry Kowal.
- Continue taking production data
- 45uA of beam current | Use ps3=0, ps5=3, ps1=-1, ps2=15, rest =-1 | One hour long runs
- Move the target to the pickup coil position and do an NMR sweep every 2/3 hours
- Return to step 1.
- Moller measurement - Thursday 11/17 (Start at 8am) (Tentative time for Mini Spin dance?)
- Optics runs with sieve in - Towards the end of this pass
- Calibration runs on H2 - We have barely enough!
- GEM luminosity study - Opportunistically | High current on C-optics | sieve out
- Polarization: 43.9% (03:30pm)
Instructions for Shift Crew
- Keep an eye out for drifting BPM positions and TS1-BBHI rate. Post a screenshot of the relevant strip-charts and GUI twice every shift.
- If beam has been down due to accelerator issues perform a checkout of the beam profile by doing harp scans and check beam centering on target using the Carbon hole target, see Simona's procedure: procedure, If in doubt about whether you should go through the beam checkout procedure, call Run Co-ordinator.
- Cross check the rate with what is expected: He3, H2
- Cross check with the expected ion chamber readings: log
- If holding field is changed for any reason, do NOT perform an NMR measurement.
Be sure to follow these procedures:
- Power down SBS, BB, and Detector HV before controlled access and allowing RadCon in the hall
- Communicate to the TO that the SBS and BB magnets will be powered down, so they can update the holding field
- See the beam recovery procedures at: HOW_TOs#Beam_Recovery_Procedure
- See the instructions to set the raster at: HOW_TOs#How_to_ask_MCC_to_set_the_desired_raster_size_of_6_mm
Experimental progress (for MCC planning meeting)
- 8.6% of accumulated charge for production data (19.9 Coulombs out of 232)
- 36% Experimental run time (52 out of 141 days)
- Long term "to do" items; not urgent, some can be done when 3He target not available
- Luminosity scans with GEMs. See procedure.
- Opportunistic BCM calibration (higher current) before pass change?
- INFN LV Scan and DAQ Stability - maybe 20-30 minutes if nothing else can run
- Work and testing on LV supplies was performed during opportunistic access. See log.
- Misc. GEM commissioning items: https://sbs.jlab.org/wiki/index.php/CommissioningPlan#BB.2FSBS_GEM_Commissioning_.28Continuation.29